# Exercising the CCPA Opt-out Right on Android: Legally Mandated but Practically Challenging

### v4

@@ 4.5. Limitations
Our study has various limitations. First, our evaluation, for example, of the applicability of the CCPA opt-out right, is based on privacy policies that app operators link from their app’s Play Store Data Safety page. It may be that a policy is not applicable to an app or the policy does not distinguish between multiple apps or the operator’s website, each of which may have different data practices. Further, by its very nature, dynamic analysis only runs code partially leaving us with an incomplete picture of apps’ behavior. Also, for our GPC and AdID opt-out evaluation we evaluated apps without user interaction, except for automatically enabling and disabling the AdID setting. Since automatically enabling or disabling the AdID may be impacted by pop-up screens or other interruptions, we manually re-analyzed apps that had unexpected AdID access and confirmed the results (Section 6.2.3). For determining opt-out effects we evaluate privacy flags, domain connections, and sharing of device identifiers. However, those are only proxies. For example, an app operator may simply decide to disregard a user’s preference despite a privacy flag being set to opt-out. Further, some apps we evaluated likely use encryption with certificate pinning. While we followed state-of-the-art methods for certificate-unpinning (Xiao et al., 2023; Kollnig et al., 2022a; Nguyen et al., 2022), we were likely not able to decrypt all encrypted traffic, which is an inherent limitation of any dynamic analysis of third-party code. For our identification of ad tracking domains we do not perform an independent evaluation and rely on previous research that led to creation of the X-Ray Tracker List (Binns et al., 2018; Kollnig, 2023). We also use a VPN to simulate a California location so that the CCPA is applicable. While we cannot exclude that some apps detected our real location behaving differently than they would for a user in California, in our GPC and AdID opt-out evaluation we did not notice such different behavior in apps with and without GPS access. Finally, we did not observe server-to-server interaction, which we consider a separate research question (Cook et al., 2020).
@@ 5. UI Opt-out Evaluation
Main Result: Only 48 of the 100 apps we evaluated implement the legally mandated CCPA opt-out setting, suggesting a broad level of non-compliance with the law.
@@ 5.2. Opt-out Effects
Figure 4. Example UIs of three apps: Leafly: Find Cannabis and CBD (Google Play, 2025a) with just a privacy policy link (top left), Vivid Seats — Event Tickets (Google Play, 2025d) with an opt-out right form (top right), and Pure Sniper: Gun Shooter Games (Google Play, 2025b) with a generic/SDK/cookie opt-out setting (bottom). Figure 1 shows an example UI of a dedicated opt-out right setting.
[…]
Example UIs of three apps.
@@
Privacy Policy 45 35%–55%
No Opt-out Setting or Privacy Policy 7 2%–12%
@@
As Table 1 shows, only 48 out of 100 apps implement the legally mandated CCPA opt-out setting. The remaining 52 apps either have just a privacy policy link (45) or no opt-out setting or privacy policy link at all (7). As an example, calculating the confidence interval for the proportion of 45/100 (45%) sample apps with just a privacy policy link, using z-scores, indicates that with 95% confidence the true proportion of apps in our app dataset with only such link is 35%–55%. Thus, even without evaluating any network traffic, our results suggest that there is a broad level of non-compliance. While the AdID setting is generally available to all apps, disabling the AdID does not appear to be effective for opting out per the CCPA (Section 6).
Turning to our network traffic analysis, some apps did not reflect the opted out state in the privacy flags they passed to third parties. Thus, even if users opt out, those apps are still selling and sharing personal information. Out of the 14 apps that included the third-party ad network Vungle (Vungle, 2025), two apps still set the ccpa status flag to opted_in in the opted out condition so that Vungle continued to collect and buy personal information. Similarly, one app also still had Google’s Restricted Data Processing (RDP) (Google, 2025) set to 0. These results show that users’ opt-outs are not always honored as app operators do not notify third parties of the opt-out so that they are in actuality not opted out. We did not see differences for connecting to ad tracking domains and disclosing device identifiers between opted out and not opted out conditions.
6. GPC and AdID Opt-out Evaluation
Main Result: While we estimate with 95% confidence that 62%–81% of apps in our app dataset must honor the CCPA’s opt-out right, many apps do not do so. For example, when sending GPC signals and disabling apps’ access to the AdID, 338 apps still had the ccpa status of the ad network Vungle set to opted in while only 26 had set it to opted out.
@@ 6.1. CCPA Opt-out Right Applicability
Figure 5. For 29/33 (88%) of the most prevalent ad tracking domains in our app dataset their privacy policies allow the “Selling” or “Sharing”/“Buying” or “Collecting” of personal information per the CCPA.
[…]
For 29/33 (88%) of the most prevalent ad tracking domains in our app dataset their privacy policies allow the “Selling” or “Sharing”/“Buying” or “Collecting” of personal information per the CCPA.
Figure 6. For 60/83 (72%) randomly sampled apps from our app dataset their privacy policies stated that the operators run a “business.”
[…]
Need to add.
Evaluating how many apps are selling or sharing personal information, we find that 1,787/1,811 (99%) of the apps in our app dataset connect to at least one of the most prevalent 29 ad tracking domains, each of which has a privacy policy stating to (1) buy or collect or (2) sell or share personal information per the CCPA (Figure 5). Further, for the random sample of 83 apps (Section 4.4), 60/83 (72%) of their policies implied that the app operators run a business per the CCPA by mentioning the applicability of the CCPA or the explicit selling or sharing of personal information (Figure 6). Thus, we estimate that 17870.72=1,2871787*0.72=1,287 apps in our app dataset are subject to the CCPA opt-out right. Calculating the confidence interval for the proportion of 60/83 (72%) sample apps, using z-scores, indicates that with 95% confidence the true proportion of app operators running a business is 63%–82%. Applying this range to the 1,787 apps that connect to at least one of the 29 most prevalent tracking domains, we arrive at an estimate of 1,126 to 1,465 (62%–81%) apps in our app dataset being subject to the CCPA opt-out right.
@@ 6.2. Opt-out Effects
We now evaluate the effectiveness of sending GPC signals and disabling the AdID for opting out on Android apps per the CCPA. Overall, we find little evidence that the apps in our app dataset honor GPC. Notably, we see no substantial impact of GPC on apps’ notification to third parties about our opt-out via the IAB’s US Privacy String (IAB, 2024) and company-specific privacy flags (Section 6.2.1). We also do not observe a decrease in the number of connections to ad tracking domains (Section 6.2.2) or the disclosure of device identifiers (Section 6.2.3).
@@ 6.2.1. Sending CCPA Privacy Flags to Ad Networks
The IAB provides an industry-wide CCPA privacy flag in the form of the US Privacy String (IAB, 2024), to be implemented as an HTTP cookie or a JavaScript property. [^23] On January 31, 2024 the US Privacy String was deprecated and integrated into the IAB’s Global Privacy Platform (GPP) that covers privacy flags for multiple US states and international jurisdictions (IAB Tech Lab, 2022a). The IAB now recommends use of GPP instead of the US Privacy String. GPP contains the usca California privacy section corresponding to the US Privacy String (IAB Tech Lab, 2022b). Since we performed our analysis before the deprecation date we focus here on the US Privacy String. Both the US Privacy String as well as its GPP equivalent can be stored on Android devices using SharedPreferences for the app context (IAB Tech Lab, 2020, 2022a). We find that when users opt out via GPC, privacy flags generally do not change to reflect users’ CCPA opt-out status. However, some apps opt users out regardless of their opt-out status.
Figure 7. Counts of US Privacy String values in the No AdID (top) and No AdID + GPC (bottom) conditions. The rightmost columns show the opt-out counts of US Privacy String values that apps in our app dataset sent to third-party domains. For example, in the No AdID condition only 7/155 apps sent a 1YY- value to inner-active.mobi indicating the opt-out status while the remaining apps sent 1YN- (146/155) and 1YNN (2/155) values indicating the opt-in status. These results did not substantially change when enabling GPC in the No AdID + GPC condition. Notably, the number of apps sending a 1YY- value stayed the same (7/160). The counts are per app and mutually exclusive for a given domain but not across domains as one app can connect to multiple domains. We only include domains with at least one US Privacy String value in our network traffic.
[…]
Counts of US Privacy String values in the No AdID (top) and No AdID + GPC (bottom) conditions.
@@
For 199/1,811 (11%) apps in our app dataset we found a US Privacy String key, us_privacy, and a corresponding US Privacy String value in their network traffic. However, the value of the US Privacy String was rarely set to opt-out, that is, rarely had a Y in its third position, which would indicate a “Yes” for a user’s status of being opted out (IAB, 2024) (Figure 7). Disabling the AdID did not lead to substantially different opt-out rates as the comparison between the AdID and No AdID conditions shows (Appendix Figure 14). Overall, the rate of 1YY- US Privacy String values, which indicate that a user’s opt-out right is honored, is low regardless of the tested condition. Our results suggest that turning on GPC, and disabling the AdId, does not help users to exercise their CCPA opt-out right as measured via the US Privacy String.
@@
0, 1 false, true opted_in, opted_out
AdID 12, 29 152, 2 339, 23
AdID + GPC 12, 30 154, 2 341, 23
No AdID 12, 31 156, 2 337, 24
No AdID + GPC 12, 31 155, 4 338, 26
Table 2. Counts of apps that share privacy flags with ad networks. For example, when sending GPC signals and disabling apps’ access to the AdID, 338 apps still had the ccpa status of the ad network Vungle set to opted in while only 26 had set it to opted out
@@
Our evaluation of company-specific privacy flags further confirms our findings. Table 2 shows counts of company-specific privacy flag values that apps in our app dataset sent with left-column values indicating opt-in status and right-column values opt-out status. Just as for the US Privacy String, company-specific privacy flags rarely changed when disabling the AdID or sending GPC signals. Generally, the apps for which we detected privacy flags with opt-out values had this status regardless of the GPC and AdID setting. It seems that apps opt out all users by default or trigger the opt-out status when they detect requests from a California IP address. It is noteworthy that Google’s rdp flag was more often set to opt-out than to opt-in, albeit, at comparatively low absolute levels. Generally it is evident that neither disabling the AdID nor sending GPC signals makes a substantial difference overall. Disabling the AdID or sending GPC signals does not have a major effect for opting out users per the CCPA as measured via company-specific privacy flags.
@@
Figure 8. Counts of apps with IsGpcEnabled set to True for each category across test conditions. An app can set IsGpcEnabled for multiple categories.
[…]
Counts of apps with IsGpcEnabled set to True for each category across test conditions.
@@ 6.2.2. Connecting to Ad Tracking Domains
Figure 9. Counts of apps in our app dataset connecting to the 29 most prevalent ad tracking domains that are subject to the CCPA opt-out right (Figure 5). There are no substantial differences between conditions.
[…]
Need to add.
Evaluating the counts of apps in our app dataset that make ad tracking domain connections, we do not find substantial differences between the four conditions (Figure 9). However, there are minor differences. Notably, the count of apps that connect to the app-measurement.com domain in the AdID condition is higher (1,047 apps) than in the No AdID + GPC condition (808 apps). The reason for this difference could be a compounding effect where disabling the AdID and sending GPC signals each leads to a slight decrease. Overall, however, the number of apps connecting to ad tracking domains in the different conditions does not indicate that sending GPC signals or opting out via disabling the AdID are effective means for exercising the CCPA opt-out right.
One other point is noteworthy. In our app dataset, all apps that connected to first-party servers also connected to third-party servers. However, 12% of apps connected exclusively to third-party servers. Those apps will receive GPC headers directly from the user’s device without reliance on redirects from a first-party server. For those apps we did not observe substantially different levels of GPC compliance. Thus, it does not appear that the communication model—receiving GPC signals via redirects from first parties vs direct GPC signals from a user’s device—plays a role for GPC compliance. This is in line with the GPC specification, which does not differentiate between the two (Zimmeck et al., 2025). Third parties are obligated to listen for GPC signals regardless from whom they receive them.
@@ 7. Discussion: Implementing Opt-out Rights on Android and Improving User Privacy
Our results suggest that many Android apps do not provide users with an effective way to exercise their CCPA opt-out right. We see these shortcomings as a platform-level challenge. To broadly improve the privacy of Android users the platform would benefit from an opt-out setting with legal meaning in place of its current AdID setting.
@@ 7.1. Android’s Opt-out Right Gap
The long-established Android AdID setting has no substantial effect on apps’ opt-out behavior (Section 6.2). Apps’ sharing and selling of personal information generally continues despite disabling the AdID. Sending GPC signals does not have substantial effect on apps’ opt-out behavior either. At the app level, our results suggest that many apps do not have a dedicated CCPA opt-out setting, and some that do are not propagating users’ opt-out preferences to third parties (Section 5.2). As a consequence, for many apps that are required to honor the CCPA opt-out right there does not appear to be any way for Android users to exercise their right.
There could be various reasons for why that is the case. It may be that some app operators are not aware of their obligation to honor users’ rights. Others may not know how they can do so and decide to put the task off for another day. In our view, our results indicate that app-level opt-out preferences are just not feasible. A similar conclusion was reached for the web (Mahoney, 2020). While a platform-level alternative would be desirable, neither GPC nor the AdID setting in its current state are working overall. Given that 63%–82% of apps in our app dataset are required to honor the CCPA opt-out right (Section 6.1), our results suggest widespread CCPA violations due to the omission of enabling users to opt out.

### v5

@@ 4.5. Limitations
Our study has various limitations. First, our evaluation of the applicability of the CCPA opt-out right, is based on privacy policies that app operators link from their app’s Play Store Data Safety page. It may be that a policy is not applicable to an app or a policy does not distinguish between multiple apps or the operator’s website, each of which may have different data practices. Further, by its very nature, dynamic analysis only runs code partially leaving us with an incomplete picture of apps’ behavior. Also, for our GPC opt-out evaluation we evaluated apps without user interaction, except for automatically enabling and disabling the AdID setting. [^23] Since automatically enabling or disabling the AdID may be impacted by pop-up screens or other interruptions, we manually re-analyzed apps that had unexpected AdID access and confirmed the results (Section 6.2.3). For determining opt-out effects we evaluate privacy flags, domain connections, and sharing of device identifiers. However, those are only proxies. For example, an app operator may simply disregard a user’s preference despite a privacy flag being set to opt-out. Our methodology does not allows us to observe server-to-server data sharing, for example, of device identifiers. Further, some apps we evaluated likely use encryption with certificate pinning. While we followed state-of-the-art methods for certificate-unpinning (Xiao et al., 2023; Kollnig et al., 2022a; Nguyen et al., 2022), we were likely not able to decrypt all encrypted traffic, which is an inherent limitation of any dynamic analysis of third-party code. For our identification of ad tracking domains we do not perform an independent evaluation and rely on previous research that led to creation of the X-Ray Tracker List (Binns et al., 2018; Kollnig, 2023). Similar to previous work (Kumar et al., 2022), we use a commercial VPN, Mullvad, to simulate a California location so that the CCPA is applicable. Apps may identify our real location, e.g., via Mullvad’s IP address or our time zone, and behave differently. Thus, our results may not be an accurate reflection of app behavior for California residents. However, we are not aware of such behavior. Notably, in our GPC opt-out evaluation we did not notice different behavior in apps with and without GPS access.
@@ 5. UI Opt-out Evaluation
Main Result: Only 48 of the 100 apps we evaluated implement the legally mandated CCPA opt-out setting, suggesting non-compliance with the law.
@@ 5.2. Opt-out Effects
Figure 4. Example UIs of three apps: Leafly: Find Cannabis and CBD (Google Play, 2025a) with just a privacy policy link (top left), Vivid Seats — Event Tickets (Google Play, 2025d) with an opt-out right form (top right), and Pure Sniper: Gun Shooter Games (Google Play, 2025b) with a generic/SDK/cookie opt-out setting (bottom). Figure 1 shows an example UI of a dedicated opt-out right setting.:
@@
Privacy Policy Link 45 35%–55%
No Opt-out Setting or Privacy Policy Link 7 2%–12%
@@
As Table 1 shows, only 48 out of 100 apps implement the legally mandated CCPA opt-out setting. The remaining 52 apps either have just a privacy policy link (45) or no opt-out setting or privacy policy link (7). As an example, calculating the confidence interval for the proportion of 45/100 (45%) apps with just a privacy policy link, using z-scores, indicates that with 95% confidence the true proportion of apps in our app dataset with only such link is 35%–55%. Thus, even without evaluating any network traffic, our results suggest the existence of non-compliance. While apps in some game categories tend to have higher percentages of non-compliant apps, non-compliance is generally observable across all app categories (Appendix Table 3).
Turning to our network traffic analysis, some apps did not reflect the opted out state in the privacy flags they passed to third parties. Thus, even if California residents opt out, those apps are still selling and sharing personal information. Out of the 14 apps that included the third-party ad network Vungle (Vungle, 2025), two apps still set the ccpa status flag to opted_in in the opted out condition so that Vungle continued to collect and buy personal information. Similarly, one app also still had Google’s Restricted Data Processing (RDP) (Google, 2025) set to 0. These results show that opt-outs are not always respected as app operators do not notify third parties of the opt-out so that they are in actuality not opted out. We did not see differences for connecting to ad tracking domains and disclosing device identifiers between opted out and not opted out conditions.
6. GPC Opt-out Evaluation
Main Result: While we estimate with 95% confidence that 62%–81% of apps in our app dataset must respect the CCPA’s opt-out right, many apps do not do so. For example, when sending GPC signals and disabling apps’ access to the AdID, 338 apps still had the ccpa status of the ad network Vungle set to opted in while only 26 had set it to opted out.
@@ 6.1. CCPA Opt-out Right Applicability
Figure 5. For 29/33 (88%) of the most prevalent ad tracking domains in our app dataset their privacy policies allow the “Selling” or “Sharing”/“Buying” or “Collecting” of personal information per the CCPA.:
Figure 6. For 60/83 (72%) randomly sampled apps from our app dataset their privacy policies stated that the operators run a “business.”:
Evaluating how many apps are selling or sharing personal information, we find that 1,787/1,811 (99%) of the apps in our app dataset connect to at least one of the most prevalent 29 ad tracking domains, each of which has a privacy policy stating to (1) buy or collect or (2) sell or share personal information per the CCPA (Figure 5). [^24] Performing the evaluation for all third-party tracking domains on the X-Ray Tracker List only increases the coverage by two apps to 1,789/1,811 (99%), which is indicative for the industry concentration in the online ad industry. Further, for the random sample of 83 apps (Section 4.4), 60/83 (72%) of their policies implied that the app operators run a business per the CCPA by mentioning the applicability of the CCPA or the explicit selling or sharing of personal information (Figure 6). Thus, we estimate that 17870.72=1,2871787*0.72=1,287 apps in our app dataset are subject to the CCPA opt-out right. Calculating the confidence interval for the proportion of 60/83 (72%) sample apps, using z-scores, indicates that with 95% confidence the true proportion of app operators running a business is 63%–82%. Applying this range to the 1,787 apps that connect to at least one of the 29 most prevalent tracking domains, we arrive at an estimate of 1,126 to 1,465 (62%–81%) apps in our app dataset being subject to the CCPA opt-out right.
@@ 6.2. Opt-out Effects
We now evaluate the effectiveness of sending GPC signals for opting out on Android apps per the CCPA and what effect, if any, disabling the AdID has. Overall, we find little evidence that the apps in our app dataset respect GPC. Notably, we see no substantial impact of GPC on apps’ notification to third parties about our opt-out via the IAB’s US Privacy String (IAB, 2024) and company-specific privacy flags (Section 6.2.1). We also do not observe a decrease in the number of connections to ad tracking domains (Section 6.2.2) or the disclosure of device identifiers (Section 6.2.3).
@@ 6.2.1. Sending CCPA Privacy Flags to Ad Networks
The IAB provides an industry-wide CCPA privacy flag in the form of the US Privacy String (IAB, 2024), to be implemented as an HTTP cookie or a JavaScript property. [^25] On January 31, 2024 the US Privacy String was deprecated and integrated into the IAB’s Global Privacy Platform (GPP) that covers privacy flags for multiple US states and international jurisdictions (IAB Tech Lab, 2022a). The IAB now recommends use of GPP instead of the US Privacy String. GPP contains the usca California privacy section corresponding to the US Privacy String (IAB Tech Lab, 2022b). Since we performed our analysis before the deprecation date we focus here on the US Privacy String. Both the US Privacy String as well as its GPP equivalent can be stored on Android devices using SharedPreferences for the app context (IAB Tech Lab, 2020, 2022a). We find that when opting out via GPC, privacy flags generally do not change to reflect the CCPA opt-out status. However, some apps opted us out regardless of whether or not we sent GPC signals or enabled or disabled the AdID.
Figure 7. Counts of US Privacy String values in the No AdID (top) and No AdID + GPC (bottom) conditions. The rightmost columns show the opt-out counts of US Privacy String values that apps in our app dataset sent to third-party domains. For example, in the No AdID condition only 7/155 apps sent a 1YY- value to inner-active.mobi indicating the opt-out status while the remaining apps sent 1YN- (146/155) and 1YNN (2/155) values indicating the opt-in status. These results did not substantially change when enabling GPC in the No AdID + GPC condition. Notably, the number of apps sending a 1YY- value stayed the same (7/160). The counts are per app and mutually exclusive for a given domain but not across domains as one app can connect to multiple domains. We only include domains with at least one US Privacy String value in our network traffic.:
@@
For 199/1,811 (11%) apps in our app dataset we found a US Privacy String key, us_privacy, and a corresponding US Privacy String value in their network traffic. However, the value of the US Privacy String was rarely set to opt-out, that is, rarely had a Y in its third position, which would indicate a “Yes” for a California resident’s status of being opted out (IAB, 2024) (Figure 7). Disabling the AdID did not lead to substantially different opt-out rates as the comparison between the AdID and No AdID conditions shows (Appendix Figure 14). Overall, the rate of 1YY- US Privacy String values, which indicate that a California resident’s opt-out right is respected, is low regardless of the tested condition. Our results suggest that turning on GPC does not help California residents to exercise their CCPA opt-out right as measured via the US Privacy String. Disabling the AdID did not change this result.
@@
0, 1, % false, true, % opted_in, opted_out, %
AdID 12, 29, 71% 152, 2, 1% 339, 23, 6%
AdID + GPC 12, 30, 71% 154, 2, 1% 341, 23, 6%
No AdID 12, 31, 72% 156, 2, 1% 337, 24, 7%
No AdID + GPC 12, 31, 72% 155, 4, 3% 338, 26, 7%
Table 2. Counts and percentages of apps that share privacy flags with ad networks. For example, when sending GPC signals and disabling apps’ access to the AdID, 338 apps still had the ccpa status of the ad network Vungle set to opted in while only 26 had set it to opted out, which corresponds to only 7% of the total apps.
@@
Our evaluation of company-specific privacy flags further confirms our findings. Table 2 shows counts and percentages of apps in our app dataset with left-column values indicating company-specific privacy flag opt-in status, middle-column values indicating opt-out status, and right-column values indicating opt-out percentages. Just as for the US Privacy String, company-specific privacy flags rarely changed when disabling the AdID or sending GPC signals. Generally, the apps for which we detected privacy flags with opt-out values had this status regardless of the GPC and AdID setting. It seems that apps opt out all California residents by default or trigger the opt-out status when they detect requests from a California IP address. It is noteworthy that Google’s rdp flag was more often set to opt-out than to opt-in, albeit, at comparatively low absolute levels. Generally it is evident that sending GPC signals does not make a substantial difference overall. Sending GPC signals does not have a major effect for opting out California residents per the CCPA as measured via company-specific privacy flags. Disabling the AdID does not lead to a different result either.
@@
Figure 8. Counts of apps with IsGpcEnabled set to True for each category across test conditions. An app can set IsGpcEnabled for multiple categories.:
@@ 6.2.2. Connecting to Ad Tracking Domains
Figure 9. Counts of apps in our app dataset connecting to the 29 most prevalent ad tracking domains that are subject to the CCPA opt-out right (Figure 5). There are no substantial differences between conditions.:
Evaluating the counts of apps in our app dataset that make ad tracking domain connections, we do not find substantial differences between the four conditions (Figure 9). However, there are minor differences. Notably, the count of apps that connect to the app-measurement.com domain in the AdID condition is higher (1,047 apps) than in the No AdID + GPC condition (808 apps). The reason for this difference could be a compounding effect where disabling the AdID and sending GPC signals each leads to a slight decrease. Overall, however, the number of apps connecting to ad tracking domains in the different conditions does not indicate that sending GPC signals is an effective means for exercising the CCPA opt-out right. Disabling the AdID did not change this result.
One other point is noteworthy. In our app dataset, all apps that connected to first-party servers also connected to third-party servers. However, 12% of apps connected exclusively to third-party servers. Those apps will receive GPC headers directly from a user’s device without reliance on redirects from a first-party server. For those apps we did not observe substantially different levels of GPC compliance. Thus, it does not appear that the communication model—receiving GPC signals via redirects from first parties vs direct GPC signals from a user’s device—plays a role for GPC compliance. This is in line with the GPC specification, which does not differentiate between the two (Zimmeck et al., 2025). Third parties are obligated to listen for GPC signals regardless from whom they receive them.
@@ 7. Discussion: Implementing Opt-out Rights on Android and Improving User Privacy
Our results suggest that many Android apps do not provide California residents with an effective way to exercise their CCPA opt-out right. We see these shortcomings as a platform-level challenge. To broadly improve the privacy of Android users the Android platform would benefit from an opt-out setting to exercise opt-out rights under applicable law in place of its current AdID setting.
@@ 7.1. Android’s Opt-out Right Gap
The long-established Android AdID setting has no substantial effect on apps’ opt-out behavior (Section 6.2). Apps’ sharing and selling of personal information also generally continues despite sending GPC signals. At the app-level, our results suggest that many apps do not have a dedicated CCPA opt-out setting, and some that do are not propagating California residents’ opt-out preferences to third parties (Section 5.2). Given that 63%–82% of apps in our app dataset are required to respect the CCPA opt-out right (Section 6.1), our results suggest that many apps violate the CCPA as they do not enable California residents to exercise their opt-out right. It is generally difficult and in many case not possible for California residents to exercise their opt-out right.
There could be various reasons for why that is the case. First, traditionally the core business model of the ad online ecosystem has been directly opposed to respecting opt-out preferences (Zuboff, 2019). Until the recent enforcement actions by the Office of the California Attorney General (State of California Department of Justice, 2022b, a, 2025) the risk of being caught for not doing so was low. Some app operators may simply accept the risk. After all, dark patterns have been presumably proliferating because companies’ A-B testing has revealed them to be profit maximizing (Luguri and Strahilevitz, 2021). The same could be true for not providing opt-out functionality.
In our view, technical and ecosystem complexities also play a major role for the lack of CCPA opt-out right compliance. Opting out is a complex coordination problem with distributed responsibilities among first parties, ad networks, CMPs, and other parties, all of which are not aligned (Hils et al., 2021). These different responsibilities become difficult to manage without addressing them at the platform-level. Thus, in our view, our results indicate that app-level opt-out preferences are just not feasible. We need a platform-level opt-out right architecture.